For a random sample Xi.X2,......, Xn, let

Find E[(S2)]
Answer #1

We have

i=1 i= 1

7t 7l 7l 7t 7t л.n.

Here we will use:

sigma^{2}=E(X_{i}^{2})-[E(X_{i})]^{2}=E(X_{i}^{2})-mu^{2}
E(X_{i}^{2})=sigma^{2}+mu^{2}

And

V(ar{X})=E(ar{X}^{2})-[E(ar{X})]^{2}=E(ar{X}^{2})-mu^{2}
E(ar{X}^{2})=rac{sigma^{2}}{n}+mu^{2}

----------

Given that

S'^{2}=rac{1}{n}left [ sum_{i=1}^{n} X^{2}_{i} -nar{X}^{2} ight ]

So,

E(S'^{2})=Eleft [rac{1}{n}left ( sum_{i=1}^{n} X^{2}_{i} -nar{X}^{2} ight ) ight ]=rac{1}{n} Eleft [left ( sum_{i=1}^{n} X^{2}_{i} -nar{X}^{2} ight ) ight ]=rac{1}{n} left [Eleft ( sum_{i=1}^{n} X^{2}_{i} ight )-Eleft (nar{X}^{2} ight ) ight ]=rac{1}{n} left [ sum_{i=1}^{n} E(X^{2}_{i}) -Eleft (nar{X}^{2} ight ) ight ]

=rac{1}{n} left [ sum_{i=1}^{n} (sigma^{2}+mu^{2}) -nmu^{2} ight ]=rac{1}{n} left [ nsigma^{2}+nmu^{2} -nleft (rac{sigma^{2}}{n}+mu^{2} ight ) ight ]

p?(n-1)

hence,

E(S'^{2})=rac{sigma^{2}(n-1)}{n}